# A Deep Reinforcement Learning-Based TCP Congestion Control Algorithm: Design, Simulation, and Evaluation

### v1

@@ II-F2 Machine Learning for TCP
Recent research has explored the potential of machine learning to improve TCP congestion control algorithms by addressing the limitations of traditional approaches like Reno and CUBIC. These traditional algorithms often struggle with dynamic network conditions due to their fixed rules and limited adaptability [41, 42]. Machine learning-based approaches have shown better performance in various network scenarios, offering improved throughput-latency balance and adaptability to changing environments [43, 44]. Reinforcement learning, in particular, has demonstrated promise in developing adaptive TCP algorithms that can learn optimal congestion control policies through interaction with the network [41, 44]. For example, QTCP, a reinforcement learning-based approach, achieved 59.5% higher throughput while maintaining low latency compared to traditional TCP [44]. These developments indicate that machine learning can provide more dynamic and adaptive congestion control algorithms, potentially overcoming the limitations of traditional methods in complex and evolving network environments [42]. On the other hand, machine learning techniques, especially RL, can easily adapt to different conditions and continuously improve their learning, thereby increasing efficiency. Machine learning techniques can analyze complex data and learn from past experiences to make informed decisions about network congestion [45]. This enables data transmission with less loss and more efficient use of network capacity.
@@ II-F3 RL-Based TCP Approaches
Recent studies have investigated various reinforcement learning (RL) approaches to improve TCP congestion control in diverse network environments. TCP-RLACC provides better data transmission and packet delivery in wireless networks by specifically adjusting the congestion window’s growth rate to network conditions [46]. TCP-Drinc utilizes deep reinforcement learning (Deep RL) to adjust the congestion window size, showing effective results in complex, dynamic networks [47]. A transfer learning-supported RL framework designed to reduce latency and increase data transfer rate in video streaming applications can distinguish between congestion-related and non-congestion-related losses [48]. Furthermore, a multi-agent RL approach used for TCP congestion control in multi-user networks performs better than traditional methods [48]. An RL-based TCP congestion avoidance mechanism in data center environments significantly reduces the impact of congestion by increasing end-to-end data transfer rate [49]. These studies clearly demonstrate the potential of RL-based approaches to improve TCP performance in various network scenarios.
@@ II-G Application Areas
TCP congestion control algorithms are traditionally used to communicate between two computers, but their application areas are not limited to these. The Internet of Things (IoT) poses challenges for reliable communication among various devices. TCP, a reliable transport protocol, needs to be adapted for IoT environments. Various studies propose solutions to improve TCP performance in IoT contexts. Verma and Kumar [50] present a congestion control policy that rapidly adjusts transmission rates according to network conditions, enhancing throughput and fairness. Hamrioui and Lorenz [51] introduce LBA-Ie, a load balancing algorithm that integrates IoT parameters into TCP’s flow control, improving QoS and energy efficiency in e-health applications. Alam [52] proposes a framework addressing reliability issues in IoT device communication, focusing on error detection, correction, and acknowledgment at the TCP layer. Gomez et al. [53] provide guidance on implementing lightweight TCP stacks for constrained node networks, discussing simplification techniques and their tradeoffs. Collectively, these studies aim to adapt TCP to the unique challenges of IoT environments, enhancing reliability, efficiency, and performance. Vehicular communication technologies, including Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Everything (V2X), aim to enhance road safety by enabling information exchange between vehicles and infrastructure [54]. TCP plays a crucial role in reliable and secure data transmission for healthcare applications. It enables efficient transfer of patient data, medical images, and real-time videos for telemedicine [55]. TCP’s congestion control mechanism is vital for managing network traffic, and advanced versions like TCP CUBIC show improved performance in healthcare networks [56]. The increasing reliance on remote healthcare services necessitates robust network infrastructure and data management solutions, including big data and cloud computing [57].

### v2

@@ II-F2 Machine Learning for TCP
Recent research has explored the potential of machine learning to improve TCP congestion control algorithms by addressing the limitations of traditional approaches like Reno and CUBIC. These traditional algorithms often struggle with dynamic network conditions due to their fixed rules and limited adaptability [Afonin2019, Wei2021]. Machine learning-based approaches have shown better performance in various network scenarios, offering improved throughput-latency balance and adaptability to changing environments [Kong2018, Li2018]. Reinforcement learning, in particular, has demonstrated promise in developing adaptive TCP algorithms that can learn optimal congestion control policies through interaction with the network [Afonin2019, Li2018]. For example, QTCP, a reinforcement learning-based approach, achieved 59.5% higher throughput while maintaining low latency compared to traditional TCP [Li2018]. These developments indicate that machine learning can provide more dynamic and adaptive congestion control algorithms, potentially overcoming the limitations of traditional methods in complex and evolving network environments [Wei2021]. On the other hand, machine learning techniques, especially RL, can easily adapt to different conditions and continuously improve their learning, thereby increasing efficiency. Machine learning techniques can analyze complex data and learn from past experiences to make informed decisions about network congestion [Sneha2020]. This enables data transmission with less loss and more efficient use of network capacity.
@@ II-F3 RL-Based TCP Approaches
Recent studies have investigated various reinforcement learning (RL) approaches to improve TCP congestion control in diverse network environments. TCP-RLACC provides better data transmission and packet delivery in wireless networks by specifically adjusting the congestion window’s growth rate to network conditions [Molia2024]. TCP-Drinc utilizes deep reinforcement learning (Deep RL) to adjust the congestion window size, showing effective results in complex, dynamic networks [Xiao2019]. A transfer learning-supported RL framework designed to reduce latency and increase data transfer rate in video streaming applications can distinguish between congestion-related and non-congestion-related losses [Kasi2021]. Furthermore, a multi-agent RL approach used for TCP congestion control in multi-user networks performs better than traditional methods [Kasi2021]. An RL-based TCP congestion avoidance mechanism in data center environments significantly reduces the impact of congestion by increasing end-to-end data transfer rate [Hassan2021]. These studies clearly demonstrate the potential of RL-based approaches to improve TCP performance in various network scenarios.
@@ II-G Application Areas
TCP congestion control algorithms are traditionally used to communicate between two computers, but their application areas are not limited to these. The Internet of Things (IoT) poses challenges for reliable communication among various devices. TCP, a reliable transport protocol, needs to be adapted for IoT environments. Various studies propose solutions to improve TCP performance in IoT contexts. Verma and Kumar [Verma2020] present a congestion control policy that rapidly adjusts transmission rates according to network conditions, enhancing throughput and fairness. Hamrioui and Lorenz [Hamrioui2017] introduce LBA-Ie, a load balancing algorithm that integrates IoT parameters into TCP’s flow control, improving QoS and energy efficiency in e-health applications. Alam [Alam2018] proposes a framework addressing reliability issues in IoT device communication, focusing on error detection, correction, and acknowledgment at the TCP layer. Gomez et al. [Gomez2021] provide guidance on implementing lightweight TCP stacks for constrained node networks, discussing simplification techniques and their tradeoffs. Collectively, these studies aim to adapt TCP to the unique challenges of IoT environments, enhancing reliability, efficiency, and performance. Vehicular communication technologies, including Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Everything (V2X), aim to enhance road safety by enabling information exchange between vehicles and infrastructure [Arena2019]. TCP plays a crucial role in reliable and secure data transmission for healthcare applications. It enables efficient transfer of patient data, medical images, and real-time videos for telemedicine [Ohashi2005]. TCP’s congestion control mechanism is vital for managing network traffic, and advanced versions like TCP CUBIC show improved performance in healthcare networks [Ahmad2018]. The increasing reliance on remote healthcare services necessitates robust network infrastructure and data management solutions, including big data and cloud computing [Kumar2021].